Two men, of mass 100 kg each, stand on a cart of mass 300 kg. The cart can roll with negligible friction along a north-south track, and everything is initially at rest. One man runs towards the north and jumps off the cart at a speed of 5.0 m/s, relative to the cart. After he has jumped, the second man runs towards the south and jumps off the cart, again with a speed of 5.0 m/s relative to the cart. Calculate the speed and direction of the cart after both men have jumped off. (6 marks)


1
Expert's answer
2022-01-18T09:35:05-0500

In this problem, we need conservation of momentum. After the first jump:


"0=mv+(m+M)u→u=-\dfrac{mv}{m+M},\\\space\\
u=-\dfrac{100·5}{100+300}=-1.25\text{ m/s (1.25 m/s South)}."

After the second jump:


"(M+m)u=m(-v)+Mv_c,\\\space\\
v_c=\dfrac{(M+m)u+mv}{M},\\\space\\
v_c=\dfrac{(300+100)(-1.25)+100·5}{300}=3.33\text{ m/s}."